# Complexity Growth and the Krylov-Wigner function

### v1

@@ Abstract
For any state in a $D$-dimensional Hilbert space with a choice of basis, one can define a discrete version of the Wigner function – a quasi-probability distribution which represents the state on a discrete phase space. The Wigner function can, in general, take on negative values, and the amount of negativity in the Wigner function has an operational meaning as a resource for quantum computation. In this note, we study the growth of Wigner negativity for a generic initial state under time evolution with chaotic Hamiltonians. We introduce the Krylov-Wigner function, i.e., the Wigner function defined with respect to the Krylov basis (with appropriate phases), and show that this choice of basis minimizes the early time growth of Wigner negativity in the large $D$ limit. We take this as evidence that the Krylov basis (with appropriate phases) is ideally suited for a dual, semi-classical description of chaotic quantum dynamics at large $D$. We also numerically study the time evolution of the Krylov-Wigner function and its negativity in random matrix theory for an initial pure state. We observe that the negativity broadly shows three phases: it rises gradually for a time of $O(\sqrt{D})$, then hits a sharp ramp and finally saturates close to its upper bound of $\sqrt{D}$.
@@ 1 Introduction
In order to further explore this proposal for “gravity as efficient classical implementation of quantum computation”, [^8] See also Czech:2017ryf ; Caputa:2018kdj for similar/related ideas. it would be helpful to study Wigner negativity and its growth in models such as DSSYK or the Saad-Shenker-Stanford matrix model Saad:2019lba . A significant obstacle is that the Wigner function and its negativity are difficult to compute analytically, but as a first step towards this goal, here we numerically study the time evolution of the Krylov-Wigner function and its negativity in Gaussian random matrix theory (GUE). We observe that the negativity broadly shows three phases (see figure 1): initially, the KW function spreads locally over phase space and its negativity shows a gradual rise. At a time $t\sim\sqrt{D}$, as the Wigner function starts to cover the entire phase space, the negativity hits a sharp ramp. After this, the Wigner negativity hits a plateau and saturates close to its upper bound of $\sqrt{D}$. On the other hand, in a generic basis, the Wigner negativity encounters a sharp ramp at $t=0$ and saturates close to its upper bound value within an $O(1)$ amount of time evolution (see figure 3). These results are consistent with the claim that the Krylov basis minimizes the early-time growth of Wigner negativity, and further suggest that this basis is useful for times not scaling exponentially in $\log\,D$.
@@ 2.1 The Wigner function
where $a,b\in\mathbb{C}$ [^9] When $a=1$ and $b=0$, this is simply the harmonic-oscillator vacuum, which has a angularly symmetric Wigner function centered at the origin. Turning on an imaginary part for $a$ causes the Wigner function to get squeezed in one direction. The real and imaginary parts of $b$ displace the center of the Wigner function. Thus, states with everywhere positive Wigner functions form a very restricted class of states in the Hilbert space. But even if we were to start with a classical/Gaussian state, the positivity of the Wigner function will not, in general, be preserved under time evolution. This only happens for at most inhomogenous, quadratic (i.e., free) Hamiltonians Soto2 . The space of such at most quadratic Hamiltonians generates a group, which we may call the “Clifford” group. The Clifford group is the subgroup of the unitary group which normalizes the HW group, or in other words, maps HW operators to HW operators. Any quantum evolution which starts from a Gaussian state and involves only Clifford gates will keep the state within the restricted class of Gaussian states. This leads to a massive reduction in the complexity of the problem: instead of having to solve for the entire wavefunction $\psi(t,q)$ for every $q$ as a function of time, we now only need to keep track of the finite set of parameters $(a(t),b(t))$. In this sense, quantum dynamics gets mapped to generalized classical dynamics for such systems, and time evolution can be efficiently simulated on a classical computer Bartlett .
@@ 2.2 The discrete Wigner function
$A(q,p)=\sum_{k,\ell=0}^{D-1}\widehat{\delta}_{2q,k+l}e^{\frac{2\pi i}{D}(k-\ell)p}\outerproduct{k}{\ell}.$ | . (19)
@@ 3.1 Krylov basis
The Krylov basis has the nice property that it minimizes the “spreading” of the wavefunction under time evolution Balasubramanian:2022tpr . This can be quantified as follows: consider the time evolved state $|\psi(t)\rangle=e^{-itH}|\psi_{0}\rangle$. For any ordered, orthonormal basis =$\mathcal{B}=\left\{|n_{\mathcal{B}}\rangle\right\}_{n=0}^{D-1}$, we can defined a measure of the spread of the state $\psi(t)$:
@@
where $c_{n}$ is any positive, monotonically increasing weight function; for e.g., a natural choice is to take $c_{n}=n$. We may think of $C_{\mathcal{B}}(t)$ as some notion of “spread complexity”, i.e., a measure of how spread out the state is in the given choice of basis. Let us say we choose $|0\rangle_{\mathcal{B}}=|\psi_{0}\rangle$, so that $C_{\mathcal{B}}(0)$ is set to zero. This does not fix anything about the remaining $(D-1)$ basis vectors of $\mathcal{B}$. But one could ask for more: one could look for a basis where the early growth of $C_{\mathcal{B}}(t)$ with respect to time is minimized. In other words, we would like a basis which successively minimizes the time derivatives $\{C^{\prime}_{\mathcal{B}}(0),C^{\prime\prime}_{\mathcal{B}}(0),\cdots\}$ and so on up to the $(D-1)$th derivative. It was shown in Balasubramanian:2022tpr that the Krylov basis minimizes all of these derivatives. In this sense, the Krylov basis quantifies the minimal and inevitable spread in the wavefunction that a state undergoes under quantum evolution. To be clear, adding individual phases to the basis vectors does not change $C_{\mathcal{B}}(t)$, so any generalized Krylov basis with individual phases also minimizes the spreading.
@@ 3.3 Minimizing the growth of negativity
It is worth taking a moment to explain what we mean by “minimizing the early time growth of the Wigner negativity”. Let $\mathcal{B}$ be some basis, and let $W_{\mathcal{B}}$ be the Wigner function defined with respect to that basis. If we wish to minimize the Wigner negativity at $t=0$, then it is clear that we must pick one of the basis vectors to be $\psi_{0}$, and without loss of generality we call this vector $|0\rangle_{\mathcal{B}}=|\psi_{0}\rangle$. With this choice, the log of the Wigner negativity at $t=0$ vanishes, and is thus also minimized. Now, consider the functions:
@@
is the negativity of the $m$th Taylor series approximation.
Definition: For two bases $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$, we say that $\mathcal{S}_{\mathcal{B}_{1}}<\mathcal{S}_{\mathcal{B}_{2}}$ if there exists some $k<D$ such that $\mathcal{N}^{(m)}_{\mathcal{B}_{1}}=\mathcal{N}^{(m)}_{\mathcal{B}_{2}}$ for all $m<k$, and $\mathcal{N}^{(k)}_{\mathcal{B}_{1}}<\mathcal{N}^{(k)}_{\mathcal{B}_{2}}$. We say that a basis $\mathcal{B}$ minimizes the growth of the Wigner negativity if for any other basis $\mathcal{B}^{\prime}$, we have $\mathcal{S}_{\mathcal{B}}\leq\mathcal{S}_{\mathcal{B}^{\prime}}$.
@@
Claim: In the large-$D$ limit, the basis $\mathcal{B}$ which minimizes the early time growth of Wigner negativity belongs to the class $\mathcal{K}$ of generalized Krylov bases.
Proof: Our proof will be inductive in nature. We will assume that for the basis to minimize the first $(d-1)$ elements of $\mathcal{S}_{\mathcal{B}}$, the basis vectors $\{|n_{\mathcal{B}}\rangle\}_{n=0}^{d-1}$ must be of the generalized Krylov form, with some particular choice of phases. Now our aim is to show that the requirement that the function $\mathcal{N}^{(d)}_{\mathcal{B}}$ be minimized forces us to pick the $d$th basis vector to be the $d$th Krylov basis vector, up to a phase. To this end, consider a basis $\mathcal{B}$ such that its first $d$ basis vectors agree with that of the Krylov basis $\mathcal{K}$, up to phases which are determined by the requirement that the first $(d-1)$ elements of $\mathcal{S}_{\mathcal{B}}$ are minimized. The time-derivatives of the Wigner function at $t=0$, with respect to the basis $\mathcal{B}$, are given by:
@@
where $\alpha_{d}\in\mathbb{R}$ and $-1\leq\alpha_{d}\leq 0$ (be choice of $\phi_{d}$), and
@@
Including terms up to $O(t^{d-1})$, the Wigner function has non-trivial support till $12Q=d-$ 1. After including the $O(t^{d})$ terms, the Wigner function in this region is exactly equal to that of the Wigner function in the generalized Krylov basis, i.e., the difference in basis vectors beyond the $(d-1)$the vector does not manifest in this region:
[…]
@@
$\sum_{p,Q\geq d/2}|W_{\mathcal{K}}^{(d)}(Q,p)|\leq\sum_{p,Q\geq d/2}|W_{\mathcal{B}}^{(d)}(Q,p)|+\frac{4|c_{d}|}{\pi}\left(d|\alpha_{d}|-\sum_{k>d}k|\alpha_{k}|\right).$ (72)
Our goal now is to show that $d|\alpha_{d}|\leq\sum_{k>d}k|\alpha_{k}|$. We know that
@@
$-2|\alpha_{d}|+|\alpha_{d}|^{2}+\sum_{k>d}|\alpha_{k}|^{2}=0.$ (74)
@@
$-|\alpha_{d}|+\sum_{k>d}|\alpha_{k}|\geq 0,$ (75)
which then implies $|\alpha_{d}|\leq\sum_{k>d}|\alpha_{k}|$. From here, it immediately follows that
$d|\alpha_{d}|\leq\sum_{k>d}k|\alpha_{k}|.$ (76)
Thus, we have shown that for any given basis $\mathcal{B}$, we can choose the phase of the $d$th generalized Krylov basis vector in a way such that we can always make the negativity of the generalized Krylov basis smaller than the given basis $\mathcal{B}$. Thus, if we pick the phase of the $d$th Krylov basis vector by simply minimizing the negativity over the choice of this phase, we are guaranteed that there exists a choice of a generalized Krylov basis which minimizes the negativity up to this order.$\Box$
@@ 4.1 Setup
Figure 1: Negativity of the Wigner function $\mathcal{N}$ as a function of the re-scaled time $\frac{t}{\sqrt{D}}$ for a randomly chosen Hamiltonian from the GUE for various values of $D$.
4.2 Spread, ramp and plateau
In figure 1, we plot the negativity of the Wigner function as a function of the re-scaled time $\frac{t}{\sqrt{D}}$. The numerical plots broadly show at least three distinct phases for the negativity: (i) a gradual early-time growth which we will call the spreading phase, (ii) a sharp ramp at a time $t\sim O(\sqrt{D})$ where the negativity grows approximately linearly with a slope (with respect to $t$) of $O(\sqrt{D})$, and (iii) saturation to a plateau at approximately $\mathcal{N}_{\text{sat.}}\sim 0.8\sqrt{D}$ at late times ($t\gg\sqrt{D}$). Note that the saturation value is close to but less than the upper bound of $\sqrt{D}$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Density plots of the Krylov-Wigner function in phase space for $D=101$ for various values of time corresponding to $\frac{t}{\sqrt{D}}=(0.125,0.5,1,1.5,2,2.25,2.5,5,1000)$.
@@ 4.2 Spread, ramp and plateau
As time increases, the Wigner function spreads in phase space (see figure 2). In the spreading phase, the Wigner negativity grows gradually with time. This growth seems to be much slower than would be the case had we picked any other computational basis. For instance, in figure 3 we compare the growth of the negativity between the Krylov basis and the original coordinate basis, i.e. the basis with respect to which the entries of the Hamiltonian are picked to be random. Note that the negativity in the coordinate basis rises sharply starting at $t=0$, with a slope of $O(\sqrt{D})$, and saturates within a time of $O(1)$. We expect this to be the behaviour in any generic basis.
In contrast, the KW negativity shows a much more gradual rise, with an $O(1)$ slope at $t=0$. To compute the slop, recall that the Wigner function can be expanded perturbatively in time as follows:
$W(t,Q,p)$ $=$ $\frac{1}{D}\sum_{k,\ell}\widehat{\delta}_{2Q,k+\ell}e^{\frac{2\pi i}{D}(k-\ell)p}\left(\bra{\ell}\rho(0)\ket{k}-it\bra{\ell}[H,\rho(0)]\ket{k}+\mathcal{O}(t^{2})\right)$
@@
Figure 3: (Left) The negativity of the Wigner function with respect to the coordinate basis in comparison with that of the Krylov-Wigner function for $D=101$. (Right) The KW negativity for small times and $D=151$. The blue curve is the numerical result, while the dashed orange curve is a straight line with slope $\frac{4}{\pi}$.
As $\frac{t}{\sqrt{D}}$ becomes $O(1)$, the Wigner function starts spreading out over the entire phase space. Up until this point, the negativity is smaller than the upper bound of $\sqrt{D}$. Once the spreading wave covers all of phase space, there is a sharp, linear rise in the negativity (see figure 1) which we call the ramp. During the ramp phase, the Wigner negativity rises almost linearly with a slope of $O(\sqrt{D})$ and hits the saturation value within an $O(1)$ amount of time. [^17] Presumably, the reason for this is that the Krylov basis minimizes the growth of the Wigner negativity up to a point, but we only have the freedom to fix $D$ basis vectors by minimization. Once all this freedom is exhausted, the negativity shoots up. After the ramp, the negativity saturates to a value of approximately $0.8\sqrt{D}$, close to its maximum possible value of $\sqrt{D}$; we call this phase the plateau. In the plateau phase, the Wigner function has spread over the entire phase space uniformly, with a large number of positive and negative pockets uniformly distributed throughout. Our numerical results are consistent with the claim that the Krylov basis minimizes the early-time growth of Wigner negativity, and further suggest that this basis is useful for times not scaling exponentially in $\log\,D$. Note that the growth of the Wigner negativity for an exponential (in $\log\,D$) amount of time followed by saturation to an exponentially large value at exponential time follows the generally expected behavior of measures of computational complexity Brown_2018 .

### v2

@@ Abstract
For any state in a $D$-dimensional Hilbert space with a choice of basis, one can define a discrete version of the Wigner function – a quasi-probability distribution which represents the state on a discrete phase space. The Wigner function can, in general, take on negative values, and the amount of negativity in the Wigner function has an operational meaning as a resource for quantum computation. In this note, we study the growth of Wigner negativity for a generic initial state under time evolution with chaotic Hamiltonians. We introduce the Krylov-Wigner function, i.e., the Wigner function defined with respect to the Krylov basis (with appropriate of phases), and show that this choice of basis minimizes the early time growth of Wigner negativity in the large $D$ limit. We take this as evidence that the Krylov basis (with appropriate phases) is ideally suited for a dual, semi-classical description of chaotic quantum dynamics at large $D$. We also numerically study the time evolution of the Krylov-Wigner function and its negativity in random matrix theory for an initial pure state. We observe that the negativity rises gradually for a time of $O(D)$ and then saturates close to its upper bound of $\sqrt{D}$.
@@ 1 Introduction
In order to further explore this proposal for “gravity as efficient classical implementation of quantum computation”, [^8] See also Czech:2017ryf; Caputa:2018kdj for similar/related ideas. it would be helpful to study Wigner negativity and its growth in models such as DSSYK or the Saad-Shenker-Stanford matrix model Saad:2019lba. A significant obstacle is that the Wigner function and its negativity are difficult to compute analytically, but as a first step towards this goal, here we numerically study the time evolution of the Krylov-Wigner function and its negativity in Gaussian random matrix theory (GUE). We observe that the negativity broadly shows three phases (see figure 1): (i) there is an initial short-lived phase which will not be explored in this paper, (ii) then there is an exponentially long phase where the KW function spreads locally over phase space and its negativity shows a gradual rise, (iii) at a time $t\sim O(D)$, as the Wigner function starts to cover the entire phase space, the negativity starts to saturate and hits a plateau close to its upper bound of $\sqrt{D}$. On the other hand, in a generic basis, the Wigner negativity encounters a sharp ramp at $t=0$ and saturates close to its upper bound value within an $O(1)$ amount of time evolution (see figure 3). These results are consistent with the claim that the Krylov basis minimizes the early-time growth of Wigner negativity, and further suggest that this basis is useful for times not scaling exponentially in $\log\,D$.
@@
Note added in v2: In v1 of this manuscript, the numerical calculations of the Wigner negativity in random matrix theory displayed a ramp at times of $O(\sqrt{D})$. This was a spurious feature caused by numerical instabilities. In v2, these issues have been fixed and the numerical plots in section 4 have been updated. In addition, the discussion below equation (60) is improved, and a minor error in equation (71) was fixed. All the main points of the original manuscript remain unaltered.
@@ 2.1 The Wigner function
where $a,b\in\mathbb{C}$. [^9] When $a=1$ and $b=0$, this is simply the harmonic-oscillator vacuum, which has a angularly symmetric Wigner function centered at the origin. Turning on an imaginary part for $a$ causes the Wigner function to get squeezed in one direction. The real and imaginary parts of $b$ displace the center of the Wigner function. Thus, states with everywhere positive Wigner functions form a very restricted class of states in the Hilbert space. But even if we were to start with a classical/Gaussian state, the positivity of the Wigner function will not, in general, be preserved under time evolution. This only happens for at most inhomogenous, quadratic (i.e., free) Hamiltonians Soto2. The space of such at most quadratic Hamiltonians generates a group, which we may call the “Clifford” group. The Clifford group is the subgroup of the unitary group which normalizes the HW group, or in other words, maps HW operators to HW operators. Any quantum evolution which starts from a Gaussian state and involves only Clifford gates will keep the state within the restricted class of Gaussian states. This leads to a massive reduction in the complexity of the problem: instead of having to solve for the entire wavefunction $\psi(t,q)$ for every $q$ as a function of time, we now only need to keep track of the finite set of parameters $(a(t),b(t))$. In this sense, quantum dynamics gets mapped to generalized classical dynamics for such systems, and time evolution can be efficiently simulated on a classical computer Bartlett.
@@ 2.2 The discrete Wigner function
$|.A(q,p)=\sum_{k,\ell=0}^{D-1}\widehat{\delta}_{2q,k+l}e^{\frac{2\pi i}{D}(k-\ell)p}\outerproduct{k}{\ell}$. (19)
@@ 3.1 Krylov basis
The Krylov basis has the nice property that it minimizes the “spreading” of the wavefunction under time evolution Balasubramanian:2022tpr. This can be quantified as follows: consider the time evolved state $|\psi(t)\rangle=e^{-itH}|\psi_{0}\rangle$. For any ordered, orthonormal basis $\mathcal{B}=\left\{|n_{\mathcal{B}}\rangle\right\}_{n=0}^{D-1}$, we can defined a measure of the spread of the state $\psi(t)$:
@@
where $c_{n}$ is any positive, monotonically increasing weight function; for e.g., a natural choice is to take $c_{n}=n$. We may think of $C_{\mathcal{B}}(t)$ as some notion of “spread complexity”, i.e., a measure of how spread out the state is in the given choice of basis. Let us say we choose $|0\rangle_{\mathcal{B}}=|\psi_{0}\rangle$, so that $C_{\mathcal{B}}(0)$ is set to zero. This does not fix anything about the remaining (D1)(D-1) basis vectors of $\mathcal{B}$. But one could ask for more: one could look for a basis where the early growth of $C_{\mathcal{B}}(t)$ with respect to time is minimized. In other words, we would like a basis which successively minimizes the time derivatives $\{C^{\prime}_{\mathcal{B}}(0),C^{\prime\prime}_{\mathcal{B}}(0),\cdots\}$ and so on up to the (D1)(D-1)th derivative. It was shown in Balasubramanian:2022tpr that the Krylov basis minimizes all of these derivatives. In this sense, the Krylov basis quantifies the minimal and inevitable spread in the wavefunction that a state undergoes under quantum evolution. To be clear, adding individual phases to the basis vectors does not change $C_{\mathcal{B}}(t)$, so any generalized Krylov basis with individual phases also minimizes the spreading.
@@ 3.3 Minimizing the growth of negativity
It is worth taking a moment to explain what we mean by “minimizing the early time growth of the Wigner negativity”. Let $\mathcal{B}$ be some basis, and let $W_{\mathcal{B}}$ be the Wigner function defined with respect to that basis. If we wish to minimize the Wigner negativity at $t=0$, then it is clear that we must pick one of the basis vectors to be $\psi_{0}$, and without loss of generality we call this vector $|0\rangle_{\mathcal{B}}=|\psi_{0}\rangle$. With this choice, the Wigner negativity at $t=0$ is minimized. Now, consider the functions:
@@
is the negativity of the $m$th Taylor series approximation. We will show that for any choice of the first basis vector in $\mathcal{B}$, there exists a phase $\phi_{1}$ such that $\mathcal{N}^{(1)}_{\mathcal{K}}\leq\mathcal{N}^{(1)}_{\mathcal{B}}$. So, one may as well take $\mathcal{B}$ to agree with $\mathcal{K}$ up to the first basis vector. Similarly, for any choice of the second basis vector of $\mathcal{B}$, there exists a phase $\phi_{2}$ such that $\mathcal{N}^{(2)}_{\mathcal{K}}\leq\mathcal{N}^{(2)}_{\mathcal{B}}$, and so on. This is essentially what we mean by the statement that the generalized Krylov basis minimizes Wigner negativity growth at early times.
@@
Claim: In the large-$D$ limit, the (generalized) Krylov basis minimizes the early time growth of Wigner negativity.
Proof: Our proof will be inductive in nature. We will assume that for the basis to minimize the first (d1)(d-1) elements of $\mathcal{S}_{\mathcal{B}}$, the basis vectors $\{|n_{\mathcal{B}}\rangle\}_{n=0}^{d-1}$ must be of the generalized Krylov form, with some particular choice of phases. Now our aim is to show that the requirement that the function $\mathcal{N}^{(d)}_{\mathcal{B}}$ be minimized forces us to pick the $d$th basis vector to be the $d$th Krylov basis vector, up to a phase. To this end, consider a basis $\mathcal{B}$ such that its first $d$ basis vectors agree with that of the Krylov basis $\mathcal{K}$, up to phases which are determined by the requirement that the first (d1)(d-1) elements of $\mathcal{S}_{\mathcal{B}}$ are minimized. The time-derivatives of the Wigner function at $t=0$, with respect to the basis $\mathcal{B}$, are given by:
@@
where $\alpha_{d}\in\mathbb{R}$ and $1\leq\alpha_{d}\leq 0$ (be choice of $\phi_{d}$), and
@@
Including terms up to $O(t^{d-1})$, the Wigner function has non-trivial support till 2Q=d12Q=d-1. After including the $O(t^{d})$ terms, the Wigner function in this region is exactly equal to that of the Wigner function in the generalized Krylov basis, i.e., the difference in basis vectors beyond the (d1)(d-1)the vector does not manifest in this region:
$\overline{W}^{(d)}_{\mathcal{B}}(Q,p)=\overline{W}^{(d)}_{\mathcal{K}}(Q,p),\;\;\;\forall\;\;2Q<d.$ (69)
@@
$\sum_{p,Q\geq d/2}|W_{\mathcal{K}}^{(d)}(Q,p)|\leq\sum_{p,Q\geq d/2}|W_{\mathcal{B}}^{(d)}(Q,p)|+\frac{4|c_{d}|}{\pi}\left(|\alpha_{d}|-\sum_{k>d}|\alpha_{k}|\right).$ (72)
Our goal now is to show that $|\alpha_{d}|\leq\sum_{k>d}|\alpha_{k}|$. We know that
@@
$2|\alpha_{d}|+|\alpha_{d}|^{2}+\sum_{k>d}|\alpha_{k}|^{2}=0.$ (74)
@@
$|\alpha_{d}|+\sum_{k>d}|\alpha_{k}|\geq 0,$ (75)
[…]
@@ 4.1 Setup
Figure 1: Negativity of the Wigner function $\mathcal{N}$ as a function of the re-scaled time $\frac{t}{D}$ for a randomly chosen Hamiltonian from the GUE for various values of $D$.
4.2 Negativity growth
In figure 1, we plot the negativity of the Wigner function as a function of the re-scaled time $\frac{t}{D}$. The numerical plots broadly show at least three distinct phases for the negativity: (i) a small early time phase which will not be explored in this paper, (ii) an exponentially long phase phase where the negativity grows gradually with time, and (iii) saturation to a plateau at approximately $\mathcal{N}_{\text{sat.}}\sim 0.8\sqrt{D}$ at late times ($t\sim O(D)$). Note that the saturation value is close to but less than the upper bound of $\sqrt{D}$.
Figure 2: Density plots of $D$ times the Krylov-Wigner function in phase space for $D=251$ for various values of time corresponding to $\frac{t}{D}=(0.0,0.05,0.15,0.2,0.25,0.3,0.35,0.45,0.55,0.65,0.8,10)$.
@@ 4.2 Negativity growth
As time increases, the Wigner function spreads in phase space (see figure 2). In the spreading phase, the Wigner negativity grows gradually with time. This growth seems to be much slower than would be the case had we picked any other computational basis. For instance, in figure 3 we compare the growth of the negativity between the Krylov basis and the original coordinate basis, i.e. the basis with respect to which the entries of the Hamiltonian are picked to be random. Note that the negativity in the coordinate basis rises sharply starting at $t=0$, and saturates to the same exponentially large value of about $0.8\sqrt{D}$ within a time of $O(1)$. We expect this to be the behaviour in any generic basis.
In contrast, the KW negativity shows a much more gradual rise, with an $O(1)$ slope at $t=0$. To compute the slope, recall that the Wigner function can be expanded perturbatively in time as follows:
[…]
@@
Figure 3: (Left) The negativity of the Wigner function with respect to the coordinate basis in comparison with that of the Krylov-Wigner function for $D=199$. (Right) The KW negativity for small times and $D=199$. The blue curve is the numerical result, while the dashed orange curve is a straight line with slope $\frac{4}{\pi}$.
As $\frac{t}{D}$ becomes $O(1)$, the Wigner function starts spreading out over the entire phase space. Up until this point, the negativity is smaller than the upper bound of $\sqrt{D}$. Once the spreading wave covers all of phase space, the negativity starts to become $O(\sqrt{D})$ and saturates to a value of approximately $0.8\sqrt{D}$, close to its maximum possible value of $\sqrt{D}$. In this plateau phase, the Wigner function has spread over the entire phase space uniformly, with a large number of positive and negative pockets uniformly distributed throughout. Our numerical results are consistent with the claim that the Krylov basis minimizes the early-time growth of Wigner negativity, and further suggest that this basis is useful for times not scaling exponentially in $\log\,D$. Note that the growth of the Wigner negativity for an exponential (in $\log\,D$) amount of time followed by saturation to an exponentially large value at exponential time follows the generally expected behavior of measures of computational complexity Brown_2018.